OpenStudy (anonymous):

The present ages of Lewis and Brown are in the ratio 1:2. Three years from now, the ages will be in the ratio 3:5. Find the present ages of Lewis and Brown. Please help! :D

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Is it hard? To me VERY hard

OpenStudy (anonymous):

let the present age of lewis=x then Brown's age=2x after 3 years= lewis's age=x+3 brown's age=2x+3 given that \[\frac{x+3}{2x+3}=\frac{3}{5}\] 5(x+3)=3(2x+3) 5x+15=6x+9 x=6 present age of lewis=6 present age of brown=2x=12

OpenStudy (anonymous):

Thank you :D

OpenStudy (anonymous):

Can you help me with my other question? It's similar to this question.